### v1

# Interaction as Explanation: A User Interaction-based Method for Explaining Image Classification Models

Abstract
In computer vision, explainable AI (xAI) methods seek to mitigate the ’black-box’ problem by making the decision-making process of deep learning models more interpretable and transparent. Traditional xAI methods concentrate on visualizing input features that influence model predictions, providing insights primarily suited for experts. In this work, we present an interaction-based xAI method that enhances user comprehension of image classification models through their interaction. Thus, we developed a web-based prototype allowing users to modify images via painting and erasing, thereby observing changes in classification results. Our approach enables users to discern critical features influencing the model’s decision-making process, aligning their mental models with the model’s logic. Experiments conducted with five images demonstrate the potential of the method to reveal feature importance through user interaction. Our work contributes a novel perspective to xAI by centering on end-user engagement and understanding, paving the way for more intuitive and accessible explainability in AI systems.
1 Introduction
Significant advancements have been achieved across multiple disciplines based on deep neural network models. Notably, models have evolved substantially in computer vision to address several tasks, such as image classification, object recognition, and semantic segmentation. From convolutional neural network (CNN) to Transformer-based models, the computer vision models have demonstrated high performance on various datasets and tasks [1, 8, 4]. The practical applications of these models are extensive and diverse, including autonomous driving, optical character recognition, medical diagnostic imaging, and commercial product recommendation.
Despite the considerable advancements of deep neural networks, their deployment in the real world is fraught with challenges. Among these, the ’black-box’ problem is a particularly prominent challenge [22, 20]. This issue refers to the opacity and interpretability of deep neural network-based models, wherein the algorithms and operational principles lack transparency, making it difficult for users to comprehend the model’s decision-making process. The black-box problem can engender unpredictability in the outcomes generated by the models, thereby potentially affecting users’ trust [7]. Therefore, it is imperative to devise solutions addressing the black-box problem to apply models across diverse sectors successfully.
Thus, prior research has proposed explainable AI (xAI) methods to mitigate the black-box problem within computer vision. The representative method is a feature-based explanation method, highlighting the input features that affect the models’ output and decision [3, 12, 10, 21, 17, 18, 13]. For instance, the Local Interpretable Model-Agnostic Explanations (LIME) technique identifies areas within input images that influence the outputs of image classification models [17]. Similarly, the gradient-weighted Class Activation Mapping (Grad-CAM) technique produces saliency maps to visualize important features based on gradients from the last layer of CNN models [18]. These techniques provide intuitive visual explanations for better model understanding. Other feature-based methods include the Testing with Concept Activation Vectors (TCAV) technique, partial dependence plots (PDP), and individual conditional expectation (ICE) methods. TCAV technique measures the impact of concepts from input images on the model’s output [10]. The partial dependence plots (PDP) and individual conditional expectation (ICE) methods illustrate how changes in inputs influence model outputs using a plot [3, 12, 9, 5]. These techniques are instrumental in highlighting the most significant features of the inputs. These xAI methods could be divided based on several criteria [12, 3, 11], such as the representation of explanations (e.g., heatmap, concepts, plots), model dependency (model-specific vs. model-agnostic), and the scope of explanation (global vs. local).
The xAI methods play a crucial role by enabling users to comprehend how models process inputs [15, 3, 12, 16]. Furthermore, depending on the nature of the model and the specifics of the explanation needed, model designers could implement the appropriate xAI technique. However, previous investigations into xAI methods have primarily focused on explanations for AI experts [11]. While end-users also require explainability and interpretability on models, current xAI methods may not fully address their requirements [14]. The end-users want to participate in the design of xAI methods [11]; however, these methods are currently developed without incorporating end-user feedback.
We propose an xAI method that enables users to understand models by interacting with image classification tasks. To facilitate the interaction-based xAI method, we have developed a web-based prototype. This prototype features a canvas that displays an image of a specific object, allowing users to paint over the image directly on the canvas. Initially, users examine the classification results of the image—including the class and confidence score— before it has been painted. Subsequently, they paint over areas that they believe do not influence the classification results. Then, the painted parts of the image are masked and removed from the image classification. After painting, they check how the classification results of the masked image differ from the initial classification results. By repeating this process, users can identify critical image features that influence the model’s input-to-output process and verify if their mental model is consistent with the model’s classification criteria. This approach can be universally applied to any classification model via our prototype.
2 Interaction-based xAI
The interaction-based xAI method enables users to understand image classification models with their painting interaction in a web-based prototype. Thus, we first implemented a web-based prototype that allows users to paint an image and provides the model’s image classification results. Then, we described the interaction scenario of users to understand the models through the prototype and painting interaction.
2.1 Prototype for Interaction-based xAI Method
Our prototype was developed as a web-based system. We utilized the React library for frontend development and the FastAPI framework for backend development. The Pytorch was used for the deep learning framework [19]. Our prototype consists of a canvas with an image of an object, a section showing the classification results along with the confidence score, buttons for painting mode and erasing mode, and a button to check the classification results.
The canvas of the prototype initially loads an image with a specific object from an image API. Users can paint unimportant features or erase painted areas on the canvas via swipe or touch interaction. Then, the painted area is masked, and the image is transmitted to the backend server for image classification.
The section displaying the classification results and confidence score presents the class name along with the confidence score from the backend server. This enables users to review the classification results of the current masked image and how much confidence the model has in these results. The confidence score is expressed as a percentage.
The buttons provide functions for the prototype. Buttons for painting and erasing modes allow users to paint and erase the image through interactions. The button for classification results masks the image’s painted area and requests the masked image’s classification results.
The backend server runs the model’s inference on the image received from the frontend and responds to the request of the frontend. Thus, we first define an image classification model. The ResNet-50 model [6], pre-trained on the ImageNet dataset [2], is employed as the model. Then, the server preprocesses the image; it is resized and normalized according to the ImageNet dataset standards. After that, the model performs the inference on the image. Finally, the server applies the softmax function on the model’s output and sends the classification result –corresponding to the image class with the highest confidence– back to the frontend.
2.2 Scenario of Interaction-based xAI Method
Through our prototype, we anticipate that users can understand the input features that affect the model’s output. Furthermore, we expect them to be able to compare their initial understanding of the model to their understanding after using the prototype. This requires iterative masked image modification based on classification results. An illustrative example of the interaction-based xAI method is depicted in Figure 1. The details are outlined below.
1.
Users first examine the image and the classification results (class name and confidence score) of the image.
2.
The users click a painting mode button and create a masked image by painting the features that they think do not affect the classification results through touch or swipe interaction.
3.
After painting, the users check the classification results of the masked image.
4.
Based on the classification results, the users modify the masked image using painting or erasing modes as needed.
5.
The users can repeat the above process to understand which input features affect the model’s output.
3 Experiments
| Image | First Interaction | Second Interaction |
| golden retriever | masking background except dog | masking all except face |
| soccer ball | masking background except ball | masking all except ball and leg |
| coffee mug | masking background except mug | masking background and handle of mug |
| bakery | masking others except for shelves | masking others except for shelves, a person, and some interior |
| cinema | masking others except for screen and stage | masking others except for screen, stage, and some seats |
We conducted experiments on five images to validate how the classification results and confidence scores are changed when our interaction-based xAI method is used to generate masked images. The five images were selected to include animals, objects, and places: a golden retriever, a soccer ball, a coffee mug, a bakery, and a cinema.
In the experiments, we painted the image by interaction with the scenario based on Section 2.2. During the experiments, we generated two masked images with two separate interactions. In the first interaction, we painted all but the most representative part of the object. For animals and objects, we painted the parts that excluded the main subjects. For places, we painted everything except the elements most representative of each place. Subsequently, based on the results of the first interaction, we modified some parts of the image during the second interaction, either by painting or erasing. More details are provided in Table 1.
3.1 Experiments Results
The masked images after the interactions and classification results (class names and confidence scores) are shown in Figure 2.
For the golden retriever image, the background was masked during the first interaction. The classification result remained as a golden retriever, but the confidence score decreased to 0.49. After the second interaction, all parts were masked except for the face of the golden retriever. The classification result was changed to labrador retriever, with a confidence score of 0.83. These interactions demonstrate that the golden retriever’s overall body and the surrounding background can influence the classification results.
For the soccer ball image, all parts except the ball were masked in the first interaction. The classification result was still a soccer ball, with a confidence score of 0.94. After the second interaction, only the soccer ball and the leg in contact with it were left unmasked. The classification result was a soccer ball, and the confidence score slightly decreased to 0.86. This shows that the object is the most critical feature for classifying the soccer ball image.
In the case of the coffee mug image, every part except for the mug itself was masked during the first interaction. The classification result was a coffee mug, with the confidence score increasing to 0.74. During the second interaction, the handle of the mug was also masked. Subsequently, the classification result changed to a pill bottle, with a confidence score of 0.44. These results indicate that the coffee mug, particularly its handle, significantly influences the model’s output.
For the bakery image, everything except the shelves was masked in the first interaction. The classification result was changed to a tobacco shop, with a confidence score of 0.08. In the second interaction, the shelves, a person, and some interior elements remained unmasked. The masked image was classified as a bakery with a low confidence score of 0.46. This indicates that the classification model can be influenced not only by the primary object (e.g., shelves) but also by the surrounding interior or environment.
In the case of the cinema image, the audience seats were masked in the first interaction, leaving only the screen and stage. The classification result was a spotlight, with a confidence score of 0.14. In the second interaction, we unmasked some of the audience seats. The classification result then shifted to a cinema, with a lower confidence score of 0.32. This demonstrates that for cinema images, not only the screen but also the seats play crucial roles in image classification.
4 Conclusions
In this paper, we present an interaction-based xAI method tailored for image classification tasks. The method facilitates the generation of masked images through user interactions, allowing users to understand the model by examining the classification results of these masked images. We conducted experiments using five image samples to observe the changes in image classification results through the proposed method. The experiments demonstrated that classification results could vary according to the masked features and the importance of features. This shows that the interaction-based xAI method can help users comprehend which features are important for classifying images with computer vision models.
However, our work has several limitations. First, the method was applied solely to a basic image classification model. Given that computer vision models encompass a wide range of tasks, it is crucial to design interactions tailored to each specific task. Furthermore, our work focused on basic interactions, but exploring other interaction types, such as pinch, zoom, and tap, could be beneficial. Moreover, future considerations might include not only the creation of masked images but also modifications to the current image. Lastly, our research did not involve a user study; we only observed the changes in classification results from sample images. To ascertain whether there is an improvement in users’ understanding of computer vision models, conducting a user study is essential. Consequently, our future work aims to design a broader range of interactions, consider various computer vision tasks and models, and conduct a rigorous user study. Despite the limitations, our method is significant for proposing an xAI method that leverages end-users’ interactions.